# Optimal and Efficient Sample Size Re-estimation: A Dynamic Cost Framework

### v1

@@ 4.1 Two sufficient conditions for strict concavity of $CP_{\theta}(z_{1},n_{2})$
$CP_{\theta}(z_{1},n_{2}(z_{1}))>\Phi\!\left(-\frac{1}{K\sqrt{\,\max n_{2}(z_{1})-n_{1}\,}}\right)$
@@
$z_{1}>\max_{n_{2}(z_{1})}\left[\frac{C_{\text{crit}}-w_{2}\!\left(K\sqrt{\,n_{2}(z_{1})-n_{1}\,}-\frac{1}{K\sqrt{\,n_{2}(z_{1})-n_{1}\,}}\right)}{w_{1}}\right],$
@@
$CP_{\theta}(z_{1},n_{2}(z_{1}))>\Phi\!\left(-\frac{1}{K\sqrt{\,\max n_{2}(z_{1})-n_{1}\,}}\right)$
@@
$z_{1}>\max_{n_{2}(z_{1})}\left[\frac{C_{\text{crit}}-w_{2}\!\left(K\sqrt{\,n_{2}(z_{1})-n_{1}\,}-\tfrac{1}{K\sqrt{\,n_{2}(z_{1})-n_{1}\,}}\right)}{w_{1}}\right].$
@@ 2. Constraint 1: $280\leq n_{2}(z_{1})\leq 420$,
On the right panel of Figure 1, the CPZ rule increases the sample size when the interim statistic $z_{1}$ falls within an interval, which they identify as $[1.187,2.338]$. By lemma 2, we can show that the conditional power is differentiable and strictly concave with respective to $n_{2}$ for any $z_{1}>0.474$. Therefore, we can deduce the implied cost function, $\gamma_{CPZ}(z_{1})$, in three regions:
@@ 5.1 Comparing $n_{2}^{JT}(z_{1})$ with $n_{2}^{LR}(z_{1})$
5.2 Effect of interim analysis timing on $n_{2}^{JT}(z_{1})$ and $n_{2}^{LR}(z_{1})$

### v2

@@ 4.1 Two sufficient conditions for strict concavity of $CP_{\theta}(z_{1},n_{2})$
$CP_{\theta}(z_{1},n_{2}(z_{1}))>\Phi\!\left(-\frac{1}{K\sqrt{\,\\ NM-n_{1}\,}}\right)$
@@
$z_{1}>\max_{n_{2}\in\mathcal{N}}\left[\frac{C_{\text{crit}}-w_{2}\!\left(K\sqrt{\,n_{2}(z_{1})-n_{1}\,}-\frac{1}{K\sqrt{\,n_{2}(z_{1})-n_{1}\,}}\right)}{w_{1}}\right],$
@@
$CP_{\theta}(z_{1},n_{2}(z_{1}))>\Phi\!\left(-\frac{1}{K\sqrt{\,\\ NM-n_{1}\,}}\right)$
@@
$z_{1}>\max_{n_{2}\in\mathcal{N}}\left[\frac{C_{\text{crit}}-w_{2}\!\left(K\sqrt{\,n_{2}(z_{1})-n_{1}\,}-\tfrac{1}{K\sqrt{\,n_{2}(z_{1})-n_{1}\,}}\right)}{w_{1}}\right].$
@@ 2. Constraint 1: $280\leq n_{2}(z_{1})\leq 420$,
On the right panel of Figure 1, the CPZ rule increases the sample size when the interim statistic $z_{1}$ falls within an interval, which they identify as $[1.187,2.338]$. By lemma 2, we can show that the conditional power is differentiable and strictly concave with respect to $n_{2}$ for any $z_{1}>0.474$. Therefore, we can deduce the implied cost function, $\gamma_{CPZ}(z_{1})$, in three regions:
@@ 5.1 Comparing $n_{2}^{JT}$ with $n_{2}^{LR}$
5.2 Effect of interim analysis timing on $n_{2}^{JT}$ and $n_{2}^{LR}$